# TD-Orch: Scalable Load-Balancing for Distributed Systems with Applications to Graph Processing

### v1

@@ 2.5. TD-Orch Evaluation
Implementation. We compare four orchestration methods: direct-push, direct-pull, sorting-based, and TD-Orch. All of them use MPI for inter-machine communication. Direct-push, direct-pull, and TD-Orch are implemented directly, while the sorting-based method is built on KaDiS (axtmann2015practical; axtmann2018lightweight; axtmann2017place; harsh2019histogram), a state-of-the-art massively parallel sorting library. All the four methods follows the orchestration interface in Figure 1.
Microbenchmark. We evaluate the orchestration frameworks on a database application that uses a distributed hash table as its storage layer. Each task fetches a data item from the hash table, performs a multiply-and-add operation, and then writes the updated value back to the hash table. Contentions in the workload are handled respectively through the four distinct orchestration strategies.
@@
We deploy the orchestration frameworks on a cluster of 16 machines, each equipped with AMD Opteron 6272 processors (64 cores, 2.1 GHz), 64 MB L3 cache, and 128 GB of memory. The machines are interconnected via a 1 Gbps network.
@@
Figure 5 shows the results of a weak-scaling experiment with varying numbers of machines. In this setup, each machine executes 100k tasks, with target data addresses drawn from a Zipf distribution (zipf2016human) under different exponents $\gamma$. More detailed results for varying $\gamma$ are provided in Table 7.
@@
Both TD-Orch and the sorting-based method are resilient to adversarial skew and exhibit near-linear weak scalability, but TD-Orch is faster. In multi-machine scenarios, TD-Orch achieves geometrically-averaged speedups of $2.68\times$, $1.48\times$ and $2.70\times$ over the other three baselines respectively.
(a) $\gamma=1.5$
(b) $\gamma=2.0$
Figure 5. Weak-scaling runtime (seconds) of hash-table microbenchmarks under Zipf with $\gamma=1.5$ and $\gamma=2.0$.
3. The TDO-GP Graph Processing System
@@
We implement a distributed graph processing system called TDO-GP on top of our TD-Orch framework to demonstrate the effectiveness of TD-Orch. Following the common practice in prior vertex-centric works (zhu2016gemini; wu2021seastar; wang2023scaleg; zhu2022vctune), TDO-GP maintains two primary storage structures: one for vertices, which hold their intermediate values and are pinned to specific machines based on a predefined schema during graph ingestion; and another for edges and their weights, which can be dynamically orchestrated by the system.
@@
Interface. Our graph processing engine is primarily executed through a combined interface of TD-Orch (Section 2) and the EdgeMap abstraction (ShunB2013; shun2015ligraplus; dhulipala2017julienne; dhulipala18scalable).
Ligra (ShunB2013; shun2015ligraplus; dhulipala2017julienne) and GBBS (dhulipala18scalable) introduce EdgeMap, which works on a single machine and takes as input a graph $G=(V,E)$, a VertexSubset $U$ (a subset of vertices), and two boolean functions, $F$ and $C$. It applies $F$ to each edge $(u,v)\in E$ iff $u\in U$ and $C(v)=true$, returning a new VertexSubset $U^{\prime}$ consisting of all vertices $v$ for which $F(u,v)$ is applied and returns true. EdgeMap supports two execution modes—sparse and dense—ensuring that it runs in $O(\sum_{u\in U}\text{deg}(u))$ work (total computation) and $O(\log n)$ span (critical path length) on arbitrary inputs.
We refer to our new interface as DistEdgeMap, which is defined in Figure 6. We treat $G$ as a directed graph. For undirected inputs, we represent each undirected edge $\{u,v\}$ as two directed edges $(u,v)$ and $(v,u)$. $U$ is a DistVertexSubset, a distributed version of VertexSubset which represents a subset of vertices across distributed machines. An execution function $f$ is applied to all edges $(u,v)\in E$ where $u\in U$, and returns a value. All returned values are aggregated by DistEdgeMap using a user-defined merge-able operator merge_value, and the result is write_back to $v$. DistEdgeMap outputs a DistVertexSubset formed up by the vertices whose write_back returns true.
@@
Figure 6. DistEdgeMap function
@@
In all these algorithms, users only need to define the function $f$ and the write_back operation using only a few lines of code; then launch DistEdgeMap without needing to know the details of the underlying execution flow. Due to space constraints, we present only the pseudocode of the BFS algorithm in Algorithm 2 as a representative example to illustrate how TDO-GP operates. Please refer to Appendix D to see how to implement much more complicated algorithms such as BC in tens of lines of codes using our interface.
Input: $G=(V,E)$, where $|V|=n$; StartVertex
Output: Dist[$1:n$]: Distrib. vertex value array
[…]
2 round=1=1
3 Function f ( (u,v): Edge ) :
4  return round IF (Dist[$u$] == round - 1) ELSE -1
5Function write_back ( v: Vertex, aggr_value ) :
6  If Dist[$v$] == -1 and aggr_value != -1 then
7     Dist[$v$] = aggr_value
8     return True
9 return False
10Function BFS ( $G$, StartVertex ) :
11  DistVertexSubset frontier = {StartVertex}
12  Dist[StartVertex] = 0
13  While frontier is not empty do
14     frontier = DistEdgeMap ($G$, frontier, f, write_back, MAX)
15     round += 11
16 return Dist[$1:n$]
17
Algorithm 2 Breadth-First Search (BFS)
3.1. Main Design
@@
Ingestion-Time Orchestration of Task-Data Flow. A distributed graph algorithm that can be expressed using DistEdgeMap can also be viewed as a task–data orchestration, whereas edges correspond to tasks and vertex intermediate values correspond to data. A key observation is that, regardless of the starting vertex, in each DistEdgeMap, the contention for data (vertex value) access induced by tasks (edges) is proportional to the degree of each vertex. Since vertex degrees are intrinsic to the structure of the graph, such contention is inevitable in all DistEdgeMap operations, regardless of the choice of starting vertex in an algorithm.
@@
Given this observation, TD-Orch can be applied once as a preprocessing phase when the graph is ingested, which is sufficient to resolve all future skew in the execution of arbitrary DistEdgeMap operations. This one-time load balancing is particularly beneficial for highly skewed graphs, such as commonly encountered power-law graphs (gonzalez2012powergraph).
@@
The preprocessing process consists of two stages. In the first stage, we begin by randomly placing each edge on an arbitrary machine and then execute one round of TD-Orch, which attempts to colocate all edges with their source vertices. Edges whose sources are low-degree vertices are now placed on the same machine. For edges with high-degree sources, a tree structure of meta-tasks (denoted as source trees) is created to represent their source-side communication. We then fix the storage location of these edges and propagate future source-vertex values through the source trees. In the second stage, another round of TD-Orch is performed, where a task is assigned to each edge to reach its destination vertices.
This again results in a tree structure of meta-tasks (denoted as destination trees), this time representing the paths along which the write_backwrite\_back values to the destination vertices will be sent in a DistEdgeMap.
Adopting such a preprocessing provides two key benefits: (i) it organizes edge storage into chunks, ensuring balanced computation load; and (ii) it establishes a tree-shaped communication framework for both get_value and write_back, ensuring balanced communication with reduced overhead.
@@
Sparse-Dense Execution. Ligra (ShunB2013) and GBBS (dhulipala18scalable) introduces single-machine sparse-dense execution. We design a new dual-mode execution flow for distributed graph processing which can switch between sparse and dense execution modes to efficiently exploit the data layouts provided by TD-Orch.
The mode is determined by the size of the current DistVertexSubset. Specifically, the sparse mode is optimized for small DistVertexSubset, while the dense mode is optimized for large ones.
@@
The choice between sparse and dense mode is dynamically determined by the size of the active frontier in the current round. Specifically, if $\Sigma_{u\in{U}}deg(u)=o(P\cdot|U|)$ where $d(\cdot)$ denotes the out-degree of inter-machine edges, the sparse mode is used; otherwise, the dense mode is employed.
This dynamic switching ensures that the execution remains both work- and communication-efficient in theory (with work and communication of $O(\Sigma_{u\in{U}}deg(u))$ as in Ligra and load-balanced due to the use of TD-Orch), while also achieving high performance in practice.
3.2. Implementation Techniques
In TDO-GP, we employ several implementation techniques, which can be grouped into three categories. Due to space limit, we summarize only the key functionalities of each group here. Detailed descriptions are provided in Appendix E.
3.2.1. Optimized Global Communication
We adopt deduplication and destination-aware broadcasting to reduce inter-machine communication, alleviating this critical bottleneck.
3.2.2. Work-Efficient Local Computation
Our design leverages ParlayLib (blelloch2020parlaylib), a state-of-the-art parallel library, to implement work-efficient computations. We enhance VertexSubset with a concurrent hash table and a concurrent bitmap, and we further develop a lightweight implementation of local EdgeMap. Moreover, DistEdgeMap accepts an optional filter_dst function, excluding vertices that are guaranteed not to be destinations in the current round.
3.2.3. Aligned Coordination between Components
We design tighter coordination between system components—between threads and scheduler, among threads to prevent cache thrashing, and between local and global execution.
3.3. Formal Guarantees
Since Theorem 2.1 applies to DistEdgeMap, all of our algorithms are work-efficient and load-balanced across machines. The theoretical bounds summarized in Table 1 are achieved when $n=\Omega(P\log^{2}{P})$. While most prior works (zhu2016gemini; mofrad2020graphite; ahmad2018la3; yan2015effective; lu2014large; yan2014pregel) mostly focus on empirically improve their systems, we place strong emphasis on graph-theoretic guarantees as a guidance for system implementation.
As demonstrated in Section 4.2, this theoretical grounding directly translates into practical performance improvements.
Table 1. Theoretical bounds of graph algorithms.
@@ 3.3. Formal Guarantees
4. Evaluation
4.1. Experimental Setup
Platform. We evaluate all methods on a cluster of 16 machines, each equipped with AMD Opteron 6272 processors (64 cores, 2.1 GHz), 64 MB L3 cache, and 128 GB of memory. The machines are interconnected via a 10 Gbps network.
Competitors. Due to their strong reported performance, we select GeminiGraph, Graphite, and LA3 as baselines. GeminiGraph (zhu2016gemini) is a computation-centric distributed framework that employs a hybrid model and cache-aware optimizations to reduce communication and improve scalability. Graphite (mofrad2020graphite) is a NUMA-aware graph analytics system using thread-based partitioning and diagonal scaling to minimize communication costs while supporting linear-algebra-based graph computations. LA3 (ahmad2018la3) is a hybrid system combining a vertex-centric interface with a linear algebra engine featuring optimizations in exploiting sparsity. Additionally, we include Pregel+ (yan2015effective; lu2014large; yan2014pregel) in one end-to-end experiment in §4.2 and GBBS (dhulipala18scalable) in the ablation study presented in §4.5.
Datasets. We evaluate all methods on Reddit (hamilton2017inductive), uk-2005 (BoVWFI; BRSLLP), Twitter-2010 (kwak2010twitter; boldi2011layered), friendster (yang2012defining), Hyperlink-2012 (meusel2015graph) and Road-USA (road).
See the first column of Table 2.
4.2. End-to-End Results
Table 2 presents a performance comparison between our proposed framework, TDO-GP, and three state-of-the-art systems: Gemini (zhu2016gemini), Graphite (mofrad2020graphite), and LA3 (ahmad2018la3), evaluated on a diverse set of graph datasets. We select the number of machines roughly proportional to the number of vertices in each dataset, except for particularly demanding datasets (e.g., Road-usa) that require excessive runtime. Overall, TDO-GP consistently achieves superior performance, outperforming the baselines in most scenarios often by substantial margins.
For social media graphs such as Twitter-2010, Friendster, and Reddit, TDO-GP achieves speedups more than $2\times$ in $19$ out of the $30$ tests we carry out. These gains stem from the ability of TDO-GP to more effectively exploit graph sparsity while minimizing inter-machine communication overhead.
@@ 4.2. End-to-End Results
It is worth noting that TDO-GP underperforms the baselines in only $2$ cases out of the $30$ experiments: Twitter-2010 PR and Road-USA PR. Detailed explanations for these exceptions are provided in Section 4.5.
We also compared TDO-GP against Pregel+ (yan2015effective; lu2014large; yan2014pregel). However, their Hadoop-based implementations are at least $5\times$ slower than all the other baselines considered here and are therefore omitted from subsequent analyses.
Table 2. End-to-end runtime comparison (in seconds) across various datasets and algorithms.
@@
diam [^2] Computing the exact diameter of large graphs is often prohibitively expensive. In our results, an asterisk (*) denotes an estimated diameter obtained using the approximation method in Ligra (ShunB2013), whereas diameters without an asterisk correspond to values reported by the dataset providers. $6^{*}$ PR 0.143 0.676 0.297 1.92

### v2

@@ 4 Case Study I: Key-Value Store
Methods. We compare four orchestration methods: TD-Orch, direct-push, direct-pull, and sorting-based. All of them use MPI for inter-machine communication. Direct-push, direct-pull, and TD-Orch are implemented directly, while the sorting-based method is built on KaDiS [axtmann2015practical, axtmann2018lightweight, axtmann2017place, harsh2019histogram], a state-of-the-art massively parallel sorting library. All four methods follow the orchestration interface in Figure˜1.
@@
Machine Details. Our experiments use a cluster of 16 machines, each equipped with AMD Opteron 6272 processors (64 cores, 2.1 GHz), 64 MB L3 cache, and 128 GB of memory. The machines are interconnected via a 10 Gbps network.
Workloads. We evaluate the orchestration frameworks on YCSB workloads [cooper2010benchmarking] A (50% reads, 50% writes), B (95% reads, 5% writes), C (read-only), and LOAD (write-only), with data access patterns following a Zipf distribution [zipf2016human], a common setting in database benchmarks (e.g., [kang2022pimtreevldb, luo2023smart]). Each task fetches an item from the key-value store, performs a multiply-and-add operation, and then optionally writes the updated value back.
@@
Results. Figure˜5 shows the results of weak-scaling experiments with $P=2,4,8,16$ machines. In this setup, each machine is initialized with 2 million tasks, with target data addresses drawn from Zipf exponents $\gamma=1.5,\ 2.0,\ 2.5$.
@@
TD-Orch is (i) resilient to adversarial skew and delivers near-linear performance under weak scaling, and (ii) significantly faster than other methods. On multi-machines, TD-Orch achieves geometrically-averaged speedups of $2.09\times$, $1.42\times$, and $2.83\times$ over the other three baselines, respectively.
Figure 5: Runtime results (in seconds) for YCSB workloads A (50% reads, 50% writes), C (read-only), and LOAD (write-only). TD-Orch scales well and outperforms prior methods. Results for YCSB-B (95% reads, 5% writes) exhibit similar trends and are omitted due to space constraints.
5 Case Study II: Graph Processing Systems
@@
As another case study, we implement a distributed graph processing system called TDO-GP on top of our TD-Orch framework to demonstrate the effectiveness of TD-Orch. Following the common practice in prior vertex-centric works [zhu2016gemini, wu2021seastar, wang2023scaleg, zhu2022vctune], TDO-GP maintains two primary storage structures: one for vertices, which hold their intermediate values and are pinned to specific machines based on a predefined schema during graph ingestion; and another for edges and their weights, which can be dynamically orchestrated by the system.
@@
Interface. Our graph processing engine is primarily executed through a combined interface of TD-Orch (Figure˜1) and the EdgeMap abstraction [ShunB2013, shun2015ligraplus, dhulipala2017julienne, dhulipala18scalable].
Ligra [ShunB2013, shun2015ligraplus, dhulipala2017julienne] and GBBS [dhulipala18scalable] introduce EdgeMap, which works on a single machine and takes as input a graph $G=(V,E)$, a VertexSubset $U$ (a subset of vertices), and two boolean functions, $F$ and $C$. It applies $F$ to each edge $(u,v)\in E$ iff $u\in U$ and $C(v)=true$, returning a new VertexSubset $U^{\prime}$ consisting of all vertices $v$ for which $F(u,v)$ is applied and returns true. EdgeMap supports two execution modes—sparse and dense—ensuring that it runs in $O(\sum_{u\in U}\text{deg}(u))$ work (total computation) and $O(\log n)$ span (critical path length) on arbitrary inputs.
We refer to our new interface as DistEdgeMap, which is defined in Figure˜6. We treat $G$ as a directed graph. For undirected inputs, we represent each undirected edge $\{u,v\}$ as two directed edges $(u,v)$ and $(v,u)$. $U$ is a DistVertexSubset, a distributed version of VertexSubset which represents a subset of vertices across distributed machines. An execution function $f$ is applied to all edges $(u,v)\in E$ where $u\in U$, and returns a value. All returned values are aggregated by DistEdgeMap using a user-defined merge-able operator merge_value, and the result is write_back to $v$. DistEdgeMap outputs a DistVertexSubset formed up by the vertices whose write_back returns true.
As shown in Figure˜6, the DistEdgeMap interface closely resembles the single-machine EdgeMap, hiding the distributed system architecture and scheduling details. This is an intentional design to simplify programming for users.
@@
Figure 6: DistEdgeMap function
@@
In all these algorithms, users only need to define the functions $f$ and write_back using only a few lines of code; then launch DistEdgeMap without needing to know the details of the underlying execution flow. Please refer to Appendix C for an implementation of a complex algorithm such as BC in fewer than 70 lines of code using our interface. This is much more concise than prior frameworks (e.g., Gemini [zhu2016gemini] requires over 400 lines).
5.1 Main Design of TDO-GP
@@
Ingestion-Time Orchestration of Task-Data Flow. A distributed graph algorithm that can be expressed using DistEdgeMap can also be viewed as a task-data orchestration, whereas edges correspond to tasks and vertex intermediate values correspond to data (see §2.1). A key observation is that, regardless of the starting vertex, in each DistEdgeMap, the contention for data (vertex value) access induced by tasks (edges) is proportional to the degree of each vertex. Since vertex degrees are intrinsic to the structure of the graph, such contention is inevitable in all DistEdgeMap operations, regardless of the choice of starting vertex in an algorithm.
@@
Given this observation, TD-Orch can be applied once as a preprocessing phase when the graph is ingested, which is sufficient to resolve all future skew in the execution of arbitrary DistEdgeMap operations. In future stages, DistEdgeMap leverages these pre-orchestrated flows rather than dynamically re-orchestrating at every stage. This one-time load balancing is particularly beneficial for highly skewed graphs, such as the commonly encountered power-law graphs [gonzalez2012powergraph].
@@
The preprocessing process consists of two stages. In the first stage, we begin by randomly placing each edge on an arbitrary machine and then execute one round of TD-Orch, which attempts to colocate all edges with their source vertices. Edges whose sources are low-degree vertices are now placed on the same machine with the vertices. For edges with high-degree sources, a tree structure in meta-tasks (denoted as source trees) is created to represent their source-side communication. We then fix the storage location of these edges and propagate future source-vertex values through the source trees. In the second stage, another round of TD-Orch is performed, where each edge, starting from its location at the end of the first stage, tries to reach its destination vertex. This again results in a tree structure in meta-tasks (called destination trees), this time representing the paths along which the values to the destination vertices will be write_backwrite\_back.
Adopting this preprocessing provides two key benefits: (i) it organizes edge storage into chunks (possibly on transit machines, rather than on source/destination machines, if highly contended), ensuring balanced computation load; and (ii) it establishes a tree-shaped communication framework for both get_value and write_back, enabling aggregation communication. This design differs from prior ghost/mirror node designs (e.g., [zhu2016gemini, gonzalez2012powergraph]), which neither use transit machines to mitigate contention nor employ tree-structured networks to improve communication costs. As a result, our design achieves better load balance and improved performance (§6).
@@
Sparse-Dense Execution. We design a dual-mode execution flow for distributed graph processing which can switch between sparse and dense modes to efficiently exploit the data layouts provided by TD-Orch. The mode is determined by the size of the current DistVertexSubset. Specifically, the sparse mode is optimized for small DistVertexSubset, while the dense mode is optimized for large ones.
@@
The choice between sparse and dense mode is dynamically determined by the size of the active frontier in the current round. Specifically, if $\Sigma_{u\in{U}}deg(u)=o(P\cdot|U|)$ where $d(\cdot)$ denotes the out-degree of inter-machine edges, the sparse mode is used; otherwise, the dense mode is used. This dynamic switching ensures that the execution remains both work- and communication-efficient in theory (with work and communication of $O(\Sigma_{u\in{U}}deg(u))$ as in Ligra and load-balanced due to TD-Orch), while also achieving high practical performance.
5.2 Implementation Techniques
In TDO-GP, we employ multiple implementation techniques, which can be grouped into three categories. Due to space limit, we summarize here the key functionalities of each group. Details are provided in Appendix D.
(T1) Optimized Global Communication. We use deduplication and destination-aware broadcasting to further reduce inter-machine communication.
(T2) Work-Efficient Local Computation. Our design leverages ParlayLib [blelloch2020parlaylib], a state-of-the-art parallel library, to implement work-efficient computations. We enhance VertexSubset with a concurrent hash table + bitmap, and further develop a lightweight local EdgeMap. Moreover, DistEdgeMap accepts an optional filter_dst function, excluding vertices that are guaranteed not to be destinations in the current round.
(T3) Aligned Coordination between Components. We design tighter coordination between system components—between threads and scheduler, among threads to prevent cache thrashing, and between local and global execution.
5.3 Formal Guarantees
Since Theorem˜1 applies to DistEdgeMap, all of our algorithms are work-efficient and load-balanced across machines. The theoretical bounds summarized in Table˜1 are achieved when $n=\Omega(P\log^{2}{P})$. While prior works [zhu2016gemini, mofrad2020graphite, ahmad2018la3, yan2015effective, lu2014large, yan2014pregel] mostly focus on empirical system improvements, we place strong emphasis on graph-theoretic guarantees as guidance for implementation. As shown in §6.2, these guarantees directly translate into practical performance improvements.
Table 1: Theoretical bounds of graph algorithms.
@@ 5.3 Formal Guarantees
6 Evaluation of Graph Processing Systems
6.1 Experimental Setup
Machine Details. We evaluate all methods on the same 16-machine cluster described in §4, each equipped with AMD Opteron 6272 processors (64 cores, 2.1 GHz), 64 MB L3 cache, 128 GB memory, and a 10 Gbps interconnect network.
Competitors. Due to their state-of-the-art performance, we select GeminiGraph [zhu2016gemini], Graphite [mofrad2020graphite], and LA3 [ahmad2018la3] as baselines. GeminiGraph is the fastest graph-algorithm-based system, and the two are the fastest linear-algebra-based systems, representing the strongest alternatives in both design families. Other earlier classic designs, such as PowerGraph [gonzalez2012powergraph], GraphX [gonzalez2014graphx], PowerLyra [chen2019powerlyra], GraphPad [anderson2016graphpad], and Giraph [apache2017giraph], are not directly compared due to their reported lower performance than that of the three baseline methods [ahmad2018la3, mofrad2020graphite, zhu2016gemini] included in our evaluation.
Additionally, we include Pregel+ [yan2015effective, lu2014large, yan2014pregel] in one end-to-end experiment in §6.2, and the state-of-the-art single-machine design GBBS [dhulipala18scalable] in the ablation study in §6.5.
Datasets. We evaluate on Reddit [hamilton2017inductive], uk-2005 [BoVWFI, BRSLLP], Twitter-2010 [kwak2010twitter, boldi2011layered], Friendster [yang2012defining], Hyperlink-2012 [meusel2015graph], and Road-USA [road], covering a wide range of dataset sizes and graph characteristics. See the first column of Table˜2.
6.2 End-to-End Results
Table˜2 presents a performance comparison between our proposed framework, TDO-GP, and three state-of-the-art systems: Gemini [zhu2016gemini], Graphite [mofrad2020graphite], and LA3 [ahmad2018la3], evaluated on various graph datasets. We select the number of machines roughly proportional to the number of vertices in each dataset, except for particularly demanding datasets (e.g., Road-usa) that require excessive runtime. Overall, TDO-GP consistently achieves superior performance, outperforming the baselines in 28 out of 30 tested scenarios, most by substantial margins.
For social media graphs such as Twitter-2010, Friendster, and Reddit, TDO-GP achieves speedups more than $2\times$ in $19$ out of the $30$ tests we carry out. These gains stem from the ability of TDO-GP to (i) more effectively exploit sparsity (sparse-dense execution in §5.1 and techniques in §5.2), and (ii) minimize inter-machine communication while achieving better load balance (ingestion-time TD-Orch in §5.1).
@@ 6.2 End-to-End Results
TDO-GP underperforms the baselines in only $2$ cases out of the $30$ experiments: Twitter PR and Road-USA PR. Detailed explanations for these exceptions will be provided in §6.5.
We also compared TDO-GP against Pregel+ [yan2015effective, lu2014large, yan2014pregel]. However, their Hadoop-based implementations are at least $5\times$ slower than all the other baselines considered here and are therefore omitted from subsequent analyses.
Table 2: End-to-end runtime comparison (in seconds) across various graph datasets and algorithms.
@@
diam [^2] Computing the exact diameter of large graphs is often prohibitively expensive. In our results, an asterisk (*) denotes an estimated diameter obtained using the approximation method in Ligra [ShunB2013], whereas diameters without an asterisk correspond to values reported by the dataset providers. $6^{*}$ PR 0.143 0.676 0.297 1.92
@@
6.3 Scalability
Strong Scaling. We run the SSSP and BC algorithms on the Twitter dataset using 1, 2, 4, 8, and 16 machines. The results in Figure˜7 demonstrate that TDO-GP exhibits near linear scalability across all machine configurations, while other methods either show poor scalability when increasing the number of machines, or exhibit poor performance.
Weak Scaling. We also evaluate weak scaling on the graph systems by varying the number of machines while keeping the graph size per machine fixed. To construct graphs on such a scale, we employ two graph generation methods: Erdős–Rényi (ER) [erd6s1960evolution], which produces unskewed graphs, and Barabási–Albert (BA) [barabasi1999emergence], which produces skewed graphs. The average number of edges per machine is fixed at 40M for a fair comparison, where in the 16-machine case the graph is of a similar size as uk-2005. For the BA model, we set the power-law exponent to $\gamma=2.2$, consistent with the measured skew observed in natural graphs reported by PowerGraph [gonzalez2012powergraph]. We run both PR and BC on $P=1,2,4,8,16$ machines. In Figure˜9, the x-axis represents the number of machines and the y-axis is the execution time. The results show that TDO-GP maintains nearly constant runtime across machine counts, demonstrating excellent weak-scaling performance, whereas prior works exhibit significant scalability degradation.